OpenStudy (anonymous):

how do i get 7 as a Pythagorean triple? please help! will medal

ganeshie8 (ganeshie8):

You may use the generating formulas for primitive pythagorean triples : \(a = x^2 - y^2\) \(b = 2xy\) \(c = x^2 + y^2\)

ganeshie8 (ganeshie8):

\[x^2-y^2 = 7\] its not so hard to solve the integers \(x,y\) because the left hand side can be factored : \[(x+y)(x-y)=7\] can be broken down as system of equations : \[x+y=7\\~\\x-y=1\] solving gives \(x=4,~y=3\)

OpenStudy (anonymous):

oh wow that's a lot

ganeshie8 (ganeshie8):

plug them in the earlier generating formulas to find \(b, c\)

OpenStudy (anonymous):

thanks ^.^